A Four bar chain ABCD has a fixed link AD = 700mm, the driving crank AB= 200mm, the driven crank CD = 400mm and the connecting link BC=900mm. Find the velocity and acceleration of Point ‘P’ midway between B and C when the angle BAD = 600 and AB rotates anticlockwise at a uniform speed of 600 rpm.



1
Expert's answer
2021-07-07T09:08:42-0400

Gives

AB=200mm

CD=400mm

BC=900mm

Angle ( BAD)=60°

Speed "(w)=2\pi n"

"w=\frac{2\pi \times600}{60}=20\pi" rad/sec

r=AB=0.2m


"v=rw=0.2\times20\times3.14=12.56m/sec"

"a_c=\frac{v^2}{r}=\frac{12.56^2}{0.2}=788.768m/sec^2"